# Bibliography.bib

### v1

@@ 6.2. Jiggling over manifolds
Suppose first that $M$ is compact and $Q$ is empty. Since the set $B(s(M),\varepsilon)\subset E$ is compact, it can be covered by finitely many fibered charts $\{\phi_{i}:U_{i}\times W_{i}\to{\mathbb{R}}^{m}\times{\mathbb{R}}^{n}\}_{i\in[I]}$, where $U_{i}$ is a chart of $M$ and $W_{i}$ is a chart of the fiber. We can assume that $\ell$ is large enough such that we can cover $K_{\ell}$ by finitely many subcomplexes $\{K^{(i)}\}_{i\in[I]}$ (not necessarily nice) with $T(\operatorname{star}^{2}K^{(i)})\subset U_{i}$.
The argument now amounts to doing induction on $i$. Starting with $s=s^{(-1)}$, we produce a sequence of sections $s^{(i)}$ via subsequent applications of Corollary 6.2. We note that linear extensions and slope-perturbations work equally well for sections as they do for maps. Concretely, in the $i$th step, using Corollary 6.2, we jiggle the section $\phi_{i}\circ s^{(i-1)}:T(|\operatorname{star}^{2}(K^{(i)})|)\to\underline{{\mathbb{R}}^{n}}$ relative to $|\cup_{j=0}^{i-1}K^{(j)}|$ and $|\operatorname{ring}(\operatorname{star}(K^{(i)}))|$. We extend the resulting section $s^{(i)}|_{|\operatorname{star}^{2}(K^{(i)})|}$ to $M$, by requiring that $s^{(i)}$ equals $s^{(i-1)}$ over $M\setminus|\operatorname{star}^{2}(K^{(i)})|$. Whenever subdivision is needed for the jiggling, we do not just subdivide $K^{(i)}$, but the entirety of $K$. In finitely many steps the proof is complete.
@@
We set again $s^{(-1)}=s$ and $K^{(-1)}=K$ and produce a sequence of sections $s^{(k)}$ that are each a piecewise solution of ${\mathcal{R}}$ on a neighborhood of $C_{i}$ with respect to the triangulation $T:|K^{(k)}|\to M$. We can assume, by possibly removing some compacts from the exhaustion, that $|\operatorname{star}^{3}(K\cap C_{k},K)|\subset C_{k+1}$.
Assume now that $s^{(k-1)}$ and $K^{(k-1)}$ have been constructed, then the strategy is illustrated in Figure 6. We define the subcomplex $L^{(k)}=K^{(k-1)}\cap(C_{k}\setminus C_{k-1})$ of $K^{(k-1)}$, and whenever we take its star or ring in the following we do so as subcomplex of $K^{(k-1)}$. Using the compact case, we then jiggle $s^{(k-1)}$ over $|\operatorname{star}^{2}(L^{(k)})|$ relative to $|\operatorname{ring}(\operatorname{star}(L^{(k)}))|$. We note that we jiggle $s^{(k-1)}$ over a manifold that generally will have a boundary, but that this does not pose any issues.
Jiggling produces a section $s^{(k)}$ over $|\operatorname{star}^{2}(L^{(k)})|$ that is piecewise smooth with respect to $T:|(\operatorname{star}^{2}(L^{(k)}))_{\ell}|\to M$ for some $\ell\in\mathbb{N}$. In particular $s^{(k)}$ is now a solution over a neighborhood of $C_{k}\setminus C_{k-1}$. We extend $s^{(k)}$ to $M$ by requiring it equals $s^{(k-1)}$ on $M\setminus|\operatorname{star}^{2}(L^{(k)})|$. We can assume that the jiggling was small enough so that $s^{(k)}$ restricted to $|\operatorname{star}^{2}(L^{(k)})|\cap C_{k-1}$ remains a solution. Then $s^{(k)}$ is a solution over a neighborhood of $C_{k}$.
We extend the $\ell$th crystalline subdivision of $\operatorname{star}^{2}(L^{(k)})$ to a subdivision $K^{(k)}$ of $K$ defined as
  • the $\ell$th crystalline subdivision of $\operatorname{star}^{2}(L^{(k)})$,
  • the 00th crystalline subdivision of $K^{(k-1)}\setminus(\operatorname{star}^{3}(L^{(k)}))$, and
@@
Since we assumed that $|\operatorname{star}^{3}(K\cap C_{k},K)|\subset C_{k+1}$, we see that $K^{(k)}$ and $K^{(k-1)}$ agree over $C_{k-2}$. In the limit we hence obtain a well-defined triangulation. ∎
@@
We now start the induction and assume that $s^{(k-1)}$ and $K^{(k-1)}$ have been constructed. We recall that the section $s^{(k)}$ is constructed such that it is piecewise smooth with respect to $T:|(\operatorname{star}^{2}(L^{(k)}))_{\ell}|\to M$, where $L^{(k)}=K^{(k)}\cap(C_{k}\setminus C_{k-1})$ is a subcomplex of $K^{(k)}$. We assume that $\ell$ is large enough that we can cone off $K^{(k-1)}_{\ell}\cap C_{k}$ within a $\delta_{k}$-neighborhood in $M$ to any finer crystalline subdivision of $K$ as in Proposition 3.12. If we do so for every $k$ (and in particular have done so for $k-2$), we can then define the subdivision $K^{(k)}$ of $K$ as the result of Proposition 3.12 such that $K^{(k)}$ equals
  • the 00th crystalline subdivision of $K^{(k-1)}\cap C_{k-2}$, and
  • the $\ell$th crystalline subdivision of $K^{(k-1)}$ in $M\setminus B(C_{k-2},\delta_{k-2})$.
@@ 7. Jiggling as an h-principle
Even though the setting of topological spaces is less natural for our purposes (see Section 7.1 for an informal discussion), we do want to make the following observation: at this point we have already proven the $\pi_{0}$-surjectivity at of the map ${\operatorname{Sol_{\operatorname{PS}}}}({\mathcal{R}})\hookrightarrow\Gamma_{\operatorname{PS}}(E)$ between topological spaces. Here we denote by $\Gamma_{\operatorname{PS}}(E)$ the space of piecewise smooth sections of $E$ and by ${\operatorname{Sol_{\operatorname{PS}}}}({\mathcal{R}})$ the subspace of piecewise smooth solutions of ${\mathcal{R}}$.
@@ 7.1. A weak homotopy equivalence between topological spaces
We now make a handful of remarks about the map $\iota:{\operatorname{Sol_{\operatorname{PS}}}}({\mathcal{R}})\hookrightarrow\Gamma_{\operatorname{PS}}(E)$. We do this to motivate the simplicial approach and pose a couple of open questions. This subsection contains no results and can otherwise be skipped.
Suppose we are interested in proving $\pi_{k}$-injectivity of $\iota$. To do so we start off with a family of sections $s:({\mathbb{D}}^{n},{\mathbb{S}}^{n-1})\to(\Gamma_{\operatorname{PS}}(E),{\operatorname{Sol_{\operatorname{PS}}}}({\mathcal{R}}))$ representing a relative homotopy class. What we would like to do is consider the induced section $\tilde{s}:M\times{\mathbb{D}}^{n}\to E$ and jiggle it. For each $t\in{\mathbb{S}}^{n-1}$, the sections $\tilde{s}_{t}:M\times\{t\}\to E$ are piecewise smooth, but with respect to different triangulations $T_{t}$ of $M\times\{t\}$. However, without further assumptions, these triangulations need not be compatible in any way. In particular, even though the sections $\tilde{s}_{t}$ are piecewise smooth, the section $\tilde{s}$ need not be. It follows that we cannot apply jiggling.
Conversely, consider a section $s:M\times{\mathbb{D}}^{n}\to E$ that is piecewise smooth with respect to a triangulation $T$ of $M\times{\mathbb{D}}^{n}$. We can then wonder whether the sections $s_{t}:M\times\{t\}\to E$ are piecewise smooth. However, the restriction of $T$ to a fiber $M\times\{t\}$ need not be a triangulation. As such, the sections $s_{t}$ may not be piecewise smooth.
These remarks suggest the following approach, which we leave for the reader to explore. Suppose that we consider a triangulation $T$ of $M\times{\mathbb{D}}^{n}$ that is in general position with respect to the fibers $M\times\{t\}$ (recall Definition 8.6). This may be enough to deduce that $T$ induces a triangulation on each fiber. Moreover, perhaps this can be used to show that every section that is piecewise with respect to $T$ defines piecewise smooth sections fiberwise. This ingredient may be enough to address the connectivity of the map $\iota$.
@@ 7.2. A weak homotopy equivalence of simplicial sets
We now turn our attention to the inclusion ${\operatorname{sSol_{\operatorname{PS}}}}({\mathcal{R}})\hookrightarrow\operatorname{Sing}(\Gamma^{0}(E))$ of the simplicial set of piecewise solutions into the simplicial set of continuous sections $E$.
We denote by $\Gamma^{0}(E)$ the space of continuous sections of $E$, endowed with the $C^{0}$-topology. We recall that $\operatorname{Sing}(\Gamma^{0}(E))$ is defined by $(\operatorname{Sing}(\Gamma^{0}(E)))_{n}=\{\Delta^{n}\to\Gamma^{0}(E)\}$, where the latter set is in bijection with $\{s:M\times\Delta^{n}\to q_{n}^{*}(E)\mid s\text{ continuous section}\}$ and $q_{n}:M\times\Delta^{n}\to M$ denotes the projection onto the first factor. The simplicial set ${\operatorname{sSol_{\operatorname{PS}}}}({\mathcal{R}})$ of piecewise solutions we define as the following.
@@
${\operatorname{sSol_{\operatorname{PS}}}}({\mathcal{R}})\in{\textbf{sSet}}$ is the simplicial set with $n$-simplices
$({\operatorname{sSol_{\operatorname{PS}}}}({\mathcal{R}}))_{n}=\left\{s:M\times\Delta^{n}\to q_{n}^{*}(E)\mid s\text{ is a piecewise smooth solution of}q_{n}^{*}({\mathcal{R}})\right\}.$
The face map $\delta_{i}:({\operatorname{sSol_{\operatorname{PS}}}}({\mathcal{R}}))_{n}\to({\operatorname{sSol_{\operatorname{PS}}}}({\mathcal{R}}))_{n-1}$ is given by restricting the solution to $M\times F_{i}$ where $F_{i}$ is the face of $\Delta^{n}$ excluding the $i$th vertex, and by identifying this face $F_{i}$ with $\Delta^{n-1}$. The degeneracy map $\sigma_{i}:({\operatorname{sSol_{\operatorname{PS}}}}({\mathcal{R}}))_{n}\to({\operatorname{sSol_{\operatorname{PS}}}}({\mathcal{R}}))_{n+1}$ is induced by the pullback along the retraction $({\operatorname{id}},r_{i}):M\times\Delta^{n+1}\to M\times\Delta^{n}$ where $r_{i}$ is the affine map collapsing the edge $\langle i,i+1\rangle$.
Here it is worth remarking that we consider triangulations of $M\times\Delta^{n}$ that are compatible with the stratification of $M\times\Delta^{n}$ induced by the stratification of $\Delta^{n}$ by its faces. Concretely, we ask that the triangulation restricts to a triangulation of each $M\times F$, where $F$ is a face. We will use the fact that any triangulation of $M\times\partial\Delta^{n}$ can be extended to a triangulation of $M\times\Delta^{n}$; this follows from the contractibility of the space of triangulations [lurie2009topics].
From the perspective of jiggling it does not matter that we now consider solutions of the pullback relation $q_{n}^{*}({\mathcal{R}})$. If ${\mathcal{R}}\subset J^{1}(E)$ is an open and fiberwise dense relation, the relation $q_{n}^{*}({\mathcal{R}})\subset J^{1}(q_{n}^{*}(E))$ is as well.
To establish a weak homotopy equivalence between ${\operatorname{sSol_{\operatorname{PS}}}}({\mathcal{R}})$ and $\operatorname{Sing}(\Gamma^{0}(E))$, we just need to unpack what this entails in terms of sections of the bundle $E$. The statement will then follow almost immediately from jiggling as in Definition 1.2. The reason is that the parametric version of jiggling, in this simplicial sense, is just a specific case of the non-parametric case.
@@
${\partial\Delta[n]}$${{\operatorname{sSol_{\operatorname{PS}}}}({\mathcal{R}})}$${\Delta[n]}$${\operatorname{Sing}(\Gamma^{0}(E)),}$
there exists a lift $\Delta[n]\to{\operatorname{sSol_{\operatorname{PS}}}}({\mathcal{R}})$ such that the upper triangle commutes and the lower triangle commutes up to homotopy.
@@
We want to remark that it is also an option to define ${\operatorname{sSol_{\operatorname{PS}}}}({\mathcal{R}})$ as consisting of pairs $(s,T)$ where additionally $s$ is piecewise smooth with respect to the triangulation $T$ of $M\times\Delta^{n}$. The face map can be defined as the restriction. One way of defining the degeneracy maps is noting that the $n$-simplices and the face maps form a semi-simplicial set that is Kan, which can hence be endowed with degeneracy maps (and degenerate simplices) which make it into a simplicial set ${\operatorname{sSol_{\operatorname{PS}}^{T}}}({\mathcal{R}})$ [rourke1971delta]. More concretely, the simplicial set can be described as the left Kan extension of the semi-simplicial set. With minor adaptations the proof of Section 1.2 follows through, proving that the inclusion ${\operatorname{sSol_{\operatorname{PS}}^{T}}}({\mathcal{R}})\hookrightarrow\operatorname{Sing}(\Gamma^{0}(E))$ is a weak homotopy equivalence. In particular this recovers that the fibers of the forgetful functor ${\operatorname{sSol_{\operatorname{PS}}^{T}}}({\mathcal{R}})\to{\operatorname{sSol_{\operatorname{PS}}}}({\mathcal{R}})$ are contractible.

### v2

@@ 6.2. Jiggling over manifolds
Suppose first that $M$ is compact and $Q$ is empty. Since the set $B(s(M),\varepsilon)\subset E$ is compact, it can be covered by finitely many fibered charts $\{\phi_{i}:U_{i}\times W_{i}\to{\mathbb{R}}^{m}\times{\mathbb{R}}^{n}\}_{i\in[I]}$, where $U_{i}$ is an open subset of $M$ and $W_{i}$ an open subset of the fiber. We can assume that $\ell$ is large enough such that we can cover $K_{\ell}$ by finitely many subcomplexes $\{K^{(i)}\}_{i\in[I]}$ (not necessarily nice) with $T(|\operatorname{star}^{2}K^{(i)}|)\subset U_{i}$.
The argument now amounts to doing induction on $i$. Starting with $s=s^{(-1)}$, we produce a sequence of sections $s^{(i)}$ via subsequent applications of Corollary 6.2. We note that linear extensions and slope-perturbations work equally well for sections as they do for maps. Concretely, in the $i$th step, using Corollary 6.2, we jiggle the section $\phi_{i}\circ s^{(i-1)}:T(|\operatorname{star}^{2}(K^{(i)})|)\to\underline{{\mathbb{R}}^{n}}$ relative to $T(|\cup_{j=0}^{i-1}K^{(j)}|)$ and $T(|\operatorname{ring}(\operatorname{star}(K^{(i)}))|)$. We extend the resulting section $s^{(i)}|_{T(|\operatorname{star}^{2}(K^{(i)})|)}$ to $M$, by requiring that $s^{(i)}$ equals $s^{(i-1)}$ over $M\setminus T(|\operatorname{star}^{2}(K^{(i)})|)$. Whenever subdivision is needed for the jiggling, we do not just subdivide $K^{(i)}$, but the entirety of $K$. In finitely many steps the proof is complete.
@@
We set again $s^{(-1)}=s$ and $K^{(-1)}=K$ and produce a sequence of sections $s^{(k)}$ that are each a piecewise solution of $\mathcal{R}$} on a neighborhood of $C_{i}$ with respect to the triangulation $T:|K^{(k)}|\to M$. We can assume, by possibly removing some compacts from the exhaustion, that $T(|\operatorname{star}^{3}(K\cap C_{k},K)|)\subset C_{k+1}$.
Assume now that $s^{(k-1)}$ and $K^{(k-1)}$ have been constructed, then the strategy is illustrated in Figure 6. We define the subcomplex $L^{(k)}$ of $K^{(k-1)}$ such that $T(|L^{(k)}|)=T(|K^{(k-1)}|)\cap(C_{k}\setminus C_{k-1})$, and whenever we take its star or ring in the following we do so as subcomplex of $K^{(k-1)}$. Using the compact case, we then jiggle $s^{(k-1)}$ over $T(|\operatorname{star}^{2}(L^{(k)})|)$ relative to $T(|\operatorname{ring}(\operatorname{star}(L^{(k)}))|)$. We note that we jiggle $s^{(k-1)}$ over a manifold that generally will have a boundary, but that this does not pose any issues.
Jiggling produces a section $s^{(k)}$ over $T(|\operatorname{star}^{2}(L^{(k)})|)$ that is piecewise smooth with respect to $T:|(\operatorname{star}^{2}(L^{(k)}))|\to M$ for some $\ell\in\mathbb{N}$. In particular $s^{(k)}$ is now a solution over a neighborhood of $C_{k}\setminus C_{k-1}$. We extend $s^{(k)}$ to $M$ by requiring it equals $s^{(k-1)}$ on $M\setminus T(|\operatorname{star}^{2}(L^{(k)})|)$. We can assume that the jiggling was small enough so that $s^{(k)}$ restricted to $T(|\operatorname{star}^{2}(L^{(k)})|)\cap C_{k-1}$ remains a solution. Then $s^{(k)}$ is a solution over a neighborhood of $C_{k}$.
[…]
  • the 0th crystalline subdivision of $K^{(k-1)}\setminus(\operatorname{star}^{3}(L^{(k)}))$, and
@@
Since we assumed that $T(|\operatorname{star}^{3}(K\cap C_{k},K)|)\subset C_{k+1}$, we see that $K^{(k)}$ and $K^{(k-1)}$ agree over $C_{k-2}$. In the limit we hence obtain a well-defined triangulation. ∎
@@
We now start the induction and assume that $s^{(k-1)}$ and $K^{(k-1)}$ have been constructed. We recall that the section $s^{(k)}$ is constructed such that it is piecewise smooth with respect to $T:|(\operatorname{star}^{2}(L^{(k)}))|\to M$, where $L^{(k)}=K^{(k)}\cap(C_{k}\setminus C_{k-1})$ is a subcomplex of $K^{(k)}$. We assume that $\ell$ is large enough that we can cone off $K^{(k-1)}_{\ell}\cap C_{k}$ within a $\delta_{k}$-neighborhood in $M$ to any finer crystalline subdivision of $K$ as in Proposition 3.14. If we do so for every $k$ (and in particular have done so for k2k-2), we can then define the subdivision $K^{(k)}$ of $K$ as the result of Proposition 3.14 such that $K^{(k)}$ equals
  • the 0th crystalline subdivision of $K^{(k-1)}\cap C_{k-2}$, and
[…]
@@ 7. Jiggling as an h-principle
Even though the setting of topological spaces is less natural for our purposes (see Section 7.1 for an informal discussion), we do want to make the following observation: at this point we have already proven the $\pi_{0}$-surjectivity of the map ${\operatorname{Sol_{\mathrm{PS}}}}({\mathcal{R}})\hookrightarrow\Gamma_{\mathrm{PS}}(E)$ between topological spaces. Here we denote by $\Gamma_{\mathrm{PS}}(E)$ the space of piecewise smooth sections of $E$ and by ${\operatorname{Sol_{\mathrm{PS}}}}({\mathcal{R}})$ the subspace of piecewise smooth solutions of $\mathcal{R}$}.
@@ 7.1. A weak homotopy equivalence between topological spaces
We now make a handful of remarks about the map $\iota:{\operatorname{Sol_{\mathrm{PS}}}}({\mathcal{R}})\hookrightarrow\Gamma_{\mathrm{PS}}(E)$. We do this to motivate the simplicial approach and pose a couple of open questions. This subsection contains no results and can otherwise be skipped.
Suppose we are interested in proving $\pi_{k}$-injectivity of $\iota$. To do so we start off with a family of sections $s:(D^{n},S^{n-1})\to(\Gamma_{\mathrm{PS}}(E),{\operatorname{Sol_{\mathrm{PS}}}}({\mathcal{R}}))$ representing a relative homotopy class. What we would like to do is consider the induced section $\tilde{s}:M\times D^{n}\to E$ and jiggle it. For each $t\in S^{n-1}$, the sections $\tilde{s}_{t}:M\times\{t\}\to E$ are piecewise smooth, but with respect to different triangulations $T_{t}$ of $M\times\{t\}$. However, without further assumptions, these triangulations need not be compatible in any way. In particular, even though the sections $\tilde{s}_{t}$ are piecewise smooth, the section $\tilde{s}$ need not be. It follows that we cannot apply jiggling.
Conversely, consider a section $s:M\times D^{n}\to E$ that is piecewise smooth with respect to a triangulation $T$ of $M\times D^{n}$. We can then wonder whether the sections $s_{t}:M\times\{t\}\to E$ are piecewise smooth. However, the restriction of $T$ to a fiber $M\times\{t\}$ need not be a triangulation. As such, the sections $s_{t}$ may not be piecewise smooth.
These remarks suggest the following approach, which we leave for the reader to explore. Suppose that we consider a triangulation $T$ of $M\times D^{n}$ that is in general position with respect to the fibers $M\times\{t\}$ (recall Definition 8.6). This may be enough to deduce that $T$ induces a triangulation on each fiber. Moreover, perhaps this can be used to show that every section that is piecewise with respect to $T$ defines piecewise smooth sections fiberwise. This ingredient may be enough to address the connectivity of the map $\iota$.
@@ 7.2. A weak homotopy equivalence of simplicial sets
We now turn our attention to the inclusion ${\operatorname{sSol_{\mathrm{PS}}}}({\mathcal{R}})\hookrightarrow\operatorname{Sing}(\Gamma^{0}(E))$ of the simplicial set of piecewise solutions into the simplicial set of continuous sections $E$.
We denote by $\Gamma^{0}(E)$ the space of continuous sections of $E$, endowed with the $C^{0}$-topology. We recall that $\operatorname{Sing}(\Gamma^{0}(E))$ is defined by $(\operatorname{Sing}(\Gamma^{0}(E)))_{n}=\{\Delta^{n}\to\Gamma^{0}(E)\}$, where the latter set is in bijection with the set $\{s:M\times\Delta^{n}\to q_{n}^{*}(E)\mid s\text{ continuous section}\}$ and $q_{n}:M\times\Delta^{n}\to M$ denotes the projection onto the first factor. The simplicial set ${\operatorname{sSol_{\mathrm{PS}}}}({\mathcal{R}})$ of piecewise solutions we define as the following.
@@
${\operatorname{sSol_{\mathrm{PS}}}}({\mathcal{R}})\in{\textbf{sSet}}$ is the simplicial set with $n$-simplices
$({\operatorname{sSol_{\mathrm{PS}}}}({\mathcal{R}}))_{n}=\left\{s:M\times\Delta^{n}\to q_{n}^{*}(E)\mid s\text{ is a piecewise smooth solution of }q_{n}^{*}({\mathcal{R}})\right\}.$
The face map $\delta_{i}:({\operatorname{sSol_{\mathrm{PS}}}}({\mathcal{R}}))_{n}\to({\operatorname{sSol_{\mathrm{PS}}}}({\mathcal{R}}))_{n-1}$ is given by restricting the solution to $M\times F_{i}$ where $F_{i}$ is the face of $\Delta^{n}$ excluding the $i$th vertex, and by identifying this face $F_{i}$ with $\Delta^{n-1}$. The degeneracy map $\sigma_{i}:({\operatorname{sSol_{\mathrm{PS}}}}({\mathcal{R}}))_{n}\to({\operatorname{sSol_{\mathrm{PS}}}}({\mathcal{R}}))_{n+1}$ is induced by the pullback along the retraction $({\operatorname{id}},r_{i}):M\times\Delta^{n+1}\to M\times\Delta^{n}$ where $r_{i}$ is the affine map collapsing the edge $\langle i,i+1\rangle$.
We make two remarks regarding the above definition. First, from the perspective of jiggling, it does not matter that we now consider solutions of the pullback relation $q_{n}^{*}({\mathcal{R}})$. If ${\mathcal{R}}\subset J^{1}(E)$ is an open and fiberwise dense relation, the relation $q_{n}^{*}({\mathcal{R}})\subset J^{1}(q_{n}^{*}(E))$ is as well. Secondly, we observe that ${\operatorname{sSol_{\mathrm{PS}}}}({\mathcal{R}})$ is a Kan complex.
Lemma 7.3.
The simplicial set ${\operatorname{sSol_{\mathrm{PS}}}}({\mathcal{R}})$ is a Kan complex.
Proof.
[…]
To establish a weak homotopy equivalence between ${\operatorname{sSol_{\mathrm{PS}}}}({\mathcal{R}})$ and $\operatorname{Sing}(\Gamma^{0}(E))$, we just need to unpack what this entails in terms of sections of the bundle $E$. The statement will then follow almost immediately from jiggling as in Definition 1.2. The reason is that the parametric version of jiggling, in this simplicial sense, is just a specific case of the non-parametric case. In the proof we will use the fact that any triangulation of $M\times\partial\Delta^{n}$ can be extended to a triangulation of $M\times\Delta^{n}$; this follows from the contractibility of the space of triangulations [21].
@@
$\partial\Delta[n]$${\operatorname{sSol_{\mathrm{PS}}}}({\mathcal{R}})$$\Delta[n]$$\operatorname{Sing}(\Gamma^{0}(E)),$}
there exists a lift $\Delta[n]\to{\operatorname{sSol_{\mathrm{PS}}}}({\mathcal{R}})$ such that the upper triangle commutes and the lower triangle commutes up to homotopy.
@@
We want to remark that it is also an option to define ${\operatorname{sSol_{\mathrm{PS}}}}({\mathcal{R}})$ as consisting of pairs $(s,T)$ where additionally $s$ is piecewise smooth with respect to the triangulation $T$ of $M\times\Delta^{n}$. The face map can be defined as the restriction. One way of defining the degeneracy maps is noting that the $n$-simplices and the face maps form a semi-simplicial set that is Kan, which can hence be endowed with degeneracy maps (and degenerate simplices) which make it into a simplicial set ${\operatorname{sSol_{\mathrm{PS}}^{T}}}({\mathcal{R}})$ [28]. More concretely, the simplicial set can be described as the left Kan extension of the semi-simplicial set. With minor adaptations the proof of Section 1.2 follows through, proving that the inclusion ${\operatorname{sSol_{\mathrm{PS}}^{T}}}({\mathcal{R}})\hookrightarrow\operatorname{Sing}(\Gamma^{0}(E))$ is a weak homotopy equivalence. In particular this recovers that the fibers of the forgetful functor ${\operatorname{sSol_{\mathrm{PS}}^{T}}}({\mathcal{R}})\to{\operatorname{sSol_{\mathrm{PS}}}}({\mathcal{R}})$ are contractible. ∎